what is the molar concentration of 5g of sodium hydroxide in 250cm^3{Na=23,H=1}
1
Expert's answer
2019-11-07T08:03:54-0500

Molar concentration is estimated from the equation:

M = n / V = m / MrV,

where M - molar concentration, n - number of moles, V - volume, Mr - molecular weight.

As Mr(NaOH) = 40 g/mol, m(NaOH) = 5 g, and V(NaOH) = 250 cm3 = 0.25 L:

M(NaOH) = 5 g / (40 g/mol × 0.25 L) = 0.5 mol/L = 0.5 M


Answer: 0.5 M